OpenStudy (anonymous):

Find all the values of t in the interval [0,2pi] cos^2(t) - sin^2(t) + sin(t) = 0

sam (.sam.):

Convert all trigs into sine \[\cos^2(t) - \sin^2(t) + \sin(t) = 0\] Using \(\cos^2(t)=1-\sin^2(t)\), \[(1-sin^2(t)) - \sin^2(t) + \sin(t) = 0\] \[-2\sin^2(t) +\sin(t) +1=0\] Factor it and solve for t

OpenStudy (anonymous):

how did you get +sin in the last part?